OpenStudy (anonymous):

sqrt 7(qrtx -7 sqrt7)

OpenStudy (anonymous):

\[\sqrt{7}(\sqrt{x}-7\sqrt{7})\]

OpenStudy (anonymous):



OpenStudy (anonymous):

@jjffkk @alexthomas @cambrige

OpenStudy (tkhunny):

Is there a problem statement? Can you do this one? a(b+c) -- Expand using teh distributive property.

OpenStudy (anonymous):

is it sqrt7x or is that just * symbol in the drawing

OpenStudy (anonymous):

its 7x

OpenStudy (anonymous):

well sqrt 7x